# LLMs as Layout designers: a spatial reasoning perspective

### v1

@@ 4.2 Qualitative Results
We present qualitative comparisons on the CGL and PKU datasets across different layout generation methods in Figure 2. DS-GAN demonstrates effective canvas utilization and produces elements with relatively consistent sizes; however, its layouts often suffer from geometric flaws such as misalignment and collisions, together with occasional underlay or text placements that interfere with salient background regions. PosterLlama achieves the most visually balanced results overall, delivering layouts that better preserve structural coherence and maintain harmony across element categories. Despite the large sizes of parameters, GPT-4o performs poorly, particularly on the CGL dataset, where frequent misalignment and weak spatial organization undermine layout quality. Qwen-instruct models show clear improvements with scaling: Qwen-3B (with LaySPA) is occasionally prone to element crowding, while Qwen-7B (with LaySPA) yields more stable and balanced arrangements, reflecting stronger spatial reasoning capabilities for layout design.
@@ 5.1 Content-aware Layout Generation
Deep generative models have long dominated content-aware layout generation, with most approaches falling into three categories: GAN-based, autoregressive, and diffusion-based methods [24, 7]. For example, CGL-GAN [3] employs a Transformer-based encoder–decoder to generate layouts from constraints, while DS-GAN [1] combines CNNs and LSTMs to condition on background images and learn distributions of design sequences for poster layouts. Diffusion-based methods, such as LayoutDM [7], leverage a discrete diffusion process with Transformer-based denoising to capture complex spatial dependencies among elements, yielding more diverse and coherent outputs than GAN-based models. To alleviate the challenge of limited training data, retrieval-augmented methods like RALF [6] enrich layout generation by fetching similar examples from external databases, though they inevitably inherit the biases and limitations of the retrieved layouts. Beyond direct generation, optimization-based refiners such as LayoutRectifier [25] focus on post-processing, introducing grid alignment strategies and differentiable containment functions to mitigate misalignment, overlap, and occlusion. Despite advances, these approaches formulate layout generation as a numerical optimization problem, which rely heavily on large-scale annotated data, risk losing semantic relationships among elements, and struggle to generalize beyond training distributions. Differing from these methods, our work frames layout design as a reinforcement learning problem, enabling exploration of diverse design possibilities while maintaining structural feasibility.
@@ 5.2 LLM-based Layout Generation
Recent advances in LLMs have prompted efforts to leverage their implicit design priors and the representational flexibility of language for layout generation. A common strategy is to serialize layouts into text-based formats such as HTML [14, 15], CSS [12], or SVG [16], thereby enabling LLMs to reason over structured specifications. For example, PosterLlama [15] formulates poster design as an HTML generation task, translating design intents into structured layouts. PosterO [16] extends this paradigm by introducing hierarchical SVG-like trees, which support more diverse design intents and accommodate non-rectangular elements. Beyond serialization, relation-aware approaches such as ReLayout [17] enrich layouts with explicit annotations (e.g., regions, saliency, margins) and apply recursive decomposition with Relation-CoT reasoning to better capture spatial dependencies and interactions. Retrieval-based methods further enhance performance by grounding generation in exemplar designs: LayoutPrompter [11] selects prompt examples with constraints resembling test cases, while CAL-RAG [13] employs an agentic framework that retrieves exemplars, proposes candidate layouts with an LLM, and iteratively refines them using feedback from a vision-language grader. However, existing LLM-based methods either rely heavily on retrieval or implicit priors, or depend on emergent capabilities such as in-context learning [12] and chain-of-thought reasoning. In contrast, our work explicitly activates and enhances LLMs’ limited spatial reasoning capabilities, guiding models to learn structural relations and multi-object spatial dependencies with hybrid geometric and aesthetic rewards. This approach enables the generation of layouts that are both structurally valid and visually harmonious, with interpretable reasoning trajectories, moving beyond reliance on priors or retrieval heuristics.
6 Conclusion
We presented LaySPA, a GRPO-based framework that equips LLM agents with explicit spatial reasoning capabilities for content-aware graphic layout generation. Within this framework, the LLM agent self-explores element placement under structural and spatial constraints, guided by a hybrid reward schema that jointly evaluates output validity, structural plausibility, and visual quality. Experimental results demonstrate that LaySPA effectively learns spatial reasoning, substantially improving layout quality, outperforming larger advanced LLMs, and achieving performance comparable to specialized state-of-the-art models. To the best of our knowledge, this is the first work to investigate repurposing LLMs as autonomous layout designers from a spatial reasoning perspective. Several promising directions remain for future work. First, LaySPA currently relies on pre-detected saliency maps, which overlook the rich visual characteristics of the canvas; incorporating visual semantics may yield stronger alignment between design intent and layout structure. Second, extending LaySPA with multi-turn reinforcement learning refinements could better simulate iterative human design behaviors. Finally, scaling LaySPA to broader applications such as user interfaces, magazines, and other structured visual media would further demonstrate its generality and practical impact.
@@ 6 Conclusion
Fig. 2: Visualization of generated layouts using different methods on (a) CGL and (b) PKU datasets. Qwen-3B and Qwen-7B results are generated with LaySPA fine-tuning.

### v3

@@ 4.2 Qualitative Results
We present qualitative comparisons on the CGL and PKU datasets across different layout generation methods in Figure 4. DS-GAN demonstrates effective canvas utilization and produces elements with relatively consistent sizes; however, its layouts often suffer from geometric flaws such as misalignment and collisions, together with occasional underlay or text placements that interfere with salient background regions. PosterLlama achieves the most visually balanced results overall, delivering layouts that better preserve structural coherence and maintain harmony across element categories. Despite the large sizes of parameters, GPT-4o performs poorly, particularly on the CGL dataset, where frequent misalignment and weak spatial organization undermine layout quality. In contrast, Qwen-7B (with LaySPA) produces more stable and balanced arrangements, demonstrating stronger spatial reasoning capabilities and overall design quality.
@@ 5.1 Content-aware Layout Generation
Deep generative models have long dominated content-aware layout generation, with most approaches falling into three categories: GAN-based, autoregressive, and diffusion-based methods [chai2023two, chai2023layoutdm]. For example, CGL-GAN [zhou2022composition] employs a Transformer-based encoder–decoder to generate layouts from constraints, while DS-GAN [hsu2023posterlayout] combines CNNs and LSTMs to condition on background images and learn distributions of design sequences for poster layouts. Diffusion-based methods, such as LayoutDM [chai2023layoutdm], leverage a discrete diffusion process with Transformer-based denoising to capture complex spatial dependencies among elements, yielding more diverse and coherent outputs than GAN-based models. To alleviate the challenge of limited training data, retrieval-augmented methods like RALF [horita2024retrieval] enrich layout generation by fetching similar examples from external databases, though they inevitably inherit the biases and limitations of the retrieved layouts. Beyond direct generation, optimization-based refiners such as LayoutRectifier [shen2025layoutrectifier] focus on post-processing, introducing grid alignment strategies and differentiable containment functions to mitigate misalignment, overlap, and occlusion. Despite advances, these approaches formulate layout generation as a numerical optimization problem, which rely heavily on large-scale annotated data, risk losing semantic relationships among elements, and struggle to generalize beyond training distributions. Differing from these methods, our work frames layout design as a reinforcement learning problem, enabling exploration of diverse design possibilities while maintaining structural feasibility.
@@ 5.2 LLM-based Layout Generation
Recent advances in LLMs have prompted efforts to leverage their implicit design priors and the representational flexibility of language for layout generation. A common strategy is to serialize layouts into text-based formats such as HTML [tang2023layoutnuwa, seol2024posterllama], CSS [feng2023layoutgpt], or SVG [hsu2025postero], thereby enabling LLMs to reason over structured specifications. For example, PosterLlama [seol2024posterllama] formulates poster design as an HTML generation task, translating design intents into structured layouts. PosterO [hsu2025postero] extends this paradigm by introducing hierarchical SVG-like trees, which support more diverse design intents and accommodate non-rectangular elements. Beyond serialization, relation-aware approaches such as ReLayout [tian2025relayout] enrich layouts with explicit annotations (e.g., regions, saliency, margins) and apply recursive decomposition with Relation-CoT reasoning to better capture spatial dependencies and interactions. Retrieval-based methods further enhance performance by grounding generation in exemplar designs: LayoutPrompter [lin2023layoutprompter] selects prompt examples with constraints resembling test cases, while CAL-RAG [forouzandehmehr2025cal] employs an agentic framework that retrieves exemplars, proposes candidate layouts with an LLM, and iteratively refines them using feedback from a vision-language grader. However, existing LLM-based methods either rely heavily on retrieval or implicit priors, or depend on emergent capabilities such as in-context learning [feng2023layoutgpt] and chain-of-thought reasoning. In contrast, our work explicitly activates and enhances LLMs’ limited spatial reasoning capabilities, guiding models to learn structural relations and multi-object spatial dependencies with hybrid geometric and aesthetic rewards. This approach enables the generation of layouts that are both structurally valid and visually harmonious, with interpretable reasoning trajectories, moving beyond reliance on priors or retrieval heuristics.
@@
Fig. 4: Visualization of generated layouts using different methods on (a) CGL and (b) PKU datasets. Qwen-7B results are generated with LaySPA fine-tuning.
6 Conclusion
We presented LaySPA, a GRPO-based framework that equips LLM agents with explicit spatial reasoning capabilities for content-aware graphic layout generation. Within this framework, the LLM agent self-explores element placement under structural and spatial constraints, guided by a hybrid reward schema that jointly evaluates output validity, structural plausibility, and visual quality. Experimental results demonstrate that LaySPA effectively learns spatial reasoning, substantially improving layout quality, outperforming larger advanced LLMs, and achieving performance comparable to specialized state-of-the-art models. To the best of our knowledge, this is the first work to investigate re-purposing LLMs as autonomous layout designers from a spatial reasoning perspective. Several promising directions remain for future work. First, LaySPA currently relies on pre-detected saliency maps, which overlook the rich visual characteristics of the canvas; incorporating visual semantics may yield stronger alignment between design intent and layout structure. Second, extending LaySPA with multi-turn reinforcement learning refinements could better simulate iterative human design behaviors. Finally, scaling LaySPA to broader applications such as user interfaces, magazines, and other structured visual media would further demonstrate its generality and practical impact.
Acknowledgement
Original images from PKU-PosterLayout dataset by Hsu et al. available at https://huggingface.co/datasets/creative-graphic-design/PKU-PosterLayout, licensed under CC BY-SA 4.0. Original images from CGL dataset by Zhou et al. available at https://huggingface.co/datasets/creative-graphic-design/CGL-Dataset, licensed under CC BY-NC-SA 4.0.